# Controlled Training Data Generation with Diffusion Models

### v1

@@ Abstract
In this work, we present a method to control a text-to-image generative model to produce training data specifically “useful” for supervised learning. Unlike previous works that employ an open-loop approach and pre-define prompts to generate new data using either a language model or human expertise, we develop an automated closed-loop system which involves two feedback mechanisms. The first mechanism uses feedback from a given supervised model and finds adversarial prompts that result in image generations that maximize the model loss. While these adversarial prompts result in diverse data informed by the model, they are not informed of the target distribution, which can be inefficient. Therefore, we introduce the second feedback mechanism that guides the generation process towards a certain target distribution. We call the method combining these two mechanisms Guided Adversarial Prompts. We perform our evaluations on different tasks, datasets and architectures, with different types of distribution shifts (spuriously correlated data, unseen domains) and demonstrate the efficiency of the proposed feedback mechanisms compared to open-loop approaches.
Figure 1: A framework to generate model- and target distribution-informed training examples. Left: An overview of how we generate training data for a given supervised model $f$ and target distribution. Suppose $g$ is a text-to-image generative model that generates images conditioned on a text prompt, $S$ and label, $y$. A supervised model trained to perform a task, e.g., image classification or depth estimation, is denoted by $f$. We aim to find prompts that would generate training data useful for $f$, and we do so via two feedback mechanisms. The first mechanism makes use of feedback from $f$. In particular, we maximize the loss on the predictions from $f$ to get Adversarial Prompts. The space of Adversarial Prompts and corresponding adversarial examples are represented by orange circles. Since there can be many adversarial examples not relevant to a particular target distribution, we introduce a second feedback mechanism that guides the prompts towards a certain target distribution (green circle). Combining the two mechanisms results in Guided Adversarial Prompts, represented by $S^{*}$. This results in generations that are both relevant to the target distribution and where $f$ does not perform well. Right: We show some examples of generations attained by our method on several datasets and tasks. Namely, Taskonomy [80] (depth estimation), iWildCam [3] (animal classification), and Waterbirds [66] (bird classification). An exemplar training image is shown in the first row and the third row illustrates examples of target distributions. The second row shows the generations from Adversarial Prompts which uses only model feedback. It finds novel “styles” that fool a given model $f$, but may not match a target distribution (third row). Guided Adversarial Prompts (fourth row) uses feedback from the target distribution and generates images that both fool the model and are similar to the target.
@@ 1 Introduction
The quality of data plays a crucial role in training generalizable deep learning models [74, 52, 21]. For a model to generalize well, its training data should be representative of the test distribution where it will be deployed. However, real world test conditions change over time, while training datasets are typically collected once and remain static due to high collection costs. We, therefore, focus on generating datasets that can adapt to novel test distributions and are more cost-efficient.
@@
Diffusion generative models [61, 33, 70, 55, 67] are trained on large-scale collections of images [69] and exhibit remarkable generalization abilities by being able to produce realistic images not seen during training. Additionally, unlike static datasets that they are trained on, these generative models allow us to adapt the generation process to produce images that follow a certain conditioning. For example, they can be conditioned on textual prompts [61] or geometric information such as depth maps [82].
@@
Recent works explore the use of diffusion models to generate training data for supervised learning with promising results [68, 18, 28]. They guide the generation process using text prompts to accomplish two goals: produce aligned image-label pairs for supervised training and adapt the generated images to a certain target distribution. These proposed methods, however, find conditioning text prompts in an open-loop way by either using a language model [18] or heuristics [68]. Therefore, they lack an automatic feedback mechanism that can refine the found text prompts to produce more curated and useful training data. Furthermore, it has been argued that being able to control the input data is a key contributor to how children are able to learn with few examples [6, 44].
@@
In this work, we propose two feedback mechanisms to find prompts for generating useful training data. The first mechanism finds prompts that result in generations that maximize the loss of a particular supervised model, thus, reflecting its failure modes. We call them Adversarial Prompts (AP). This mechanism ensures that we find not only novel prompts, which may produce images that the model already performs well on, but adversarial prompts that produce images with high loss, and, thus, useful for improving the model [14] (see exemplar generations for AP in Figs. 1, 3 and 5)
@@
A given model can perform poorly on multiple distribution shifts, and traversing all of them with adversarial optimization to adapt it to a specific target distribution can be inefficient (e.g., see the difference between AP generations and the illustrated target distributions in Fig. 1-right). Therefore, we introduce an additional target-informed feedback mechanism that finds prompts that generate images similar to those from the target distribution we want to adapt to. To implement this, we assume either access to a textual description of the target distribution or a few sampled (unlabeled) images from it. We then optimize a similarity metric between CLIP [57] embeddings of the generated examples and the target description. We call these prompts Guided Adversarial Prompts (GAP). Compare the columns Adversarial Prompts and Guided Adversarial Prompts in Figs. 1, 3, 4 and 5 to see the effect of CLIP guidance in steering the generations towards a specific target distribution.
@@
We demonstrate the effectiveness of our method on different tasks (image classification, depth estimation), datasets with distribution shifts (Waterbirds [66], iWildCam [4, 41], Common Corruptions [29], 3D Common Corruptions [37]), and architectures (convolutional and transformer) with supportive results.
@@ 2 Related Work
Open-loop data generation methods use pre-defined controls to guide the generative process and produce novel training examples. One line of work uses GANs [35, 8, 59] and pre-define a perturbation in their latent space to generate novel examples. More recent works adopt text-to-image diffusion models and use pre-defined prompt templates [68, 77, 27] or use a language model to generative variations of a given prompt [77]. These methods require anticipating the kind of data that will be seen at test-time when defining the prompts. On the other hand, our CLIP guidance mechanism allows us to generate images similar to the target distribution. [18] also approach this problem by using a captioning and language model to summarize a target distribution shift into a text prompt. However, this summarization process is not informed of the generations, and, thus, does not guarantee that the text prompt will guide the generation process to images related to the target distribution. Finally, these methods are not model-informed and do not necessarily generate images useful for training a given model.
@@
Closed-loop data generation methods guide the generation process via an automatic feedback mechanism. They control the latent space of GANs [5] or VAEs [76] models, NeRF [15], or the space of hand-crafted augmentations [10] to generate data that maximizes the loss of the network on the generated data. Similarly, [36] uses an SVM to identify the failure modes of a given model and uses this information to generate training data with a diffusion model. Our method employs a similar adversarial formulation (in conjunction with target distribution guidance) but performs the optimization in the text prompt space of recently developed diffusion models.
@@
“Shallow” data augmentation techniques apply simple hand-crafted transformations to training images to increase data diversity and improve the model’s generalization. Examples of such transformations are color jitter, random crop, and flipping, etc. To produce more diverse augmentations, methods like RandAugment [11] and AugMix [30] combine multiple of such simple transformations, and Mixup [81] and CutMix [78] methods use transformations that can combine multiple images. AutoAugment [10] and adversarial training [49] build a closed system to tune the parameters of the applied augmentations but are inherently limited by the expressiveness of the simple transformations. In contrast, our method uses expressive diffusion models, which results in images that are more diverse and realistic than those produced by “shallow” augmentations.
@@
Controlling diffusion models. Methods like ControlNet [82] and T2I-Adapter [54] adapt a pre-trained diffusion model to allow for additional conditioning e.g., edge, segmentation, and depth maps. We employ these models for generation as it allows us to generate paired data for different tasks, given the labels from an existing dataset. Editing methods aim to modify a given image, either via the prompt [32], masks [9], instructions [7] or inversion of the latent space [53, 34]. In contrast, personalization methods aim to adapt diffusion models to a given concept e.g., an object, individual, or style. Popular examples include textual inversion [22] and DreamBooth [64], which aim to find a token to represent a concept given several images of that concept. The former freezes the diffusion model, while the latter fine-tunes it. Extensions of these works learn to represent multiple concepts [1, 24]. In our work, we adopt an approach similar to textual inversion to steer the diffusion model, but our method can also be used with other controlling mechanisms.
@@ 3.1 Preliminaries
Problem Formulation. We consider the problem of supervised learning, where a model $f:\mathcal{X}\to\mathcal{Y}$ learns a mapping from the image space $\mathcal{X}$, to a target space $\mathcal{Y}$, e.g., a depth estimation or semantic classification problem. The model $f$ is trained using training dataset $\mathcal{D}_{\text{train}}$ and tested on a new set $\mathcal{D}_{\text{test}}$ that exhibits a distribution shift w.r.t. the training data. Our goal is to generate additional synthetic training data $\mathcal{D}_{\text{syn}}$ to adapt the model and improve its performance under the distribution shift. To apply the target-informed feedback mechanism described in Sec. 3.3, we assume access to some information about the test distribution, either from text descriptions or a few samples of unlabeled images.
Text-to-image Diffusion Models. We use the Stable Diffusion [62] text-to-image diffusion model as the basis for our generator $g$. Given a textual prompt $c$, Stable Diffusion is capable of synthesizing realistic images following the textual conditioning. However, in general, for a given task, e.g., depth estimation, a textual prompt alone may not be sufficient for controlling the generation well enough to produce aligned image-label examples.
Generating aligned training examples. We employ the following two approaches to condition the generative model $g$ on the label $y$ and sample aligned training examples $(\tilde{x},y)$. For the depth estimation task, we use the ControlNet [83] model which extends the conditioning mechanisms of the Stable Diffusion to accept various spatial modalities, e.g., depth maps, segmentation masks, or edges. Specifically, we use ControlNet v1.0 with depth conditioning [^1] https://github.com/lllyasviel/ControlNet. For semantic classification tasks, we utilize the foreground object masks and use an in-painting technique proposed in [47] that preserves the masked region throughout the denoising process, essentially keeping it intact. These mechanisms provide us with a generative model conditioned both on a text prompt $c$ and label $y$. We denote the resulting distribution modeled by this generative model as $g(y,c)$.
@@ 3.2 Model-Informed Generation with Adversarial Prompt Optimization
Our first feedback mechanism aims at generating training examples that reflect the failure modes of a given model $f$. An automatic way to do so is via adversarial optimization, which finds the “worst case” failure modes of $f$. More precisely, we find a text prompt $c$ that generates images $\tilde{x}\sim g(y,c)$ that maximize the supervised loss $\mathcal{L}(f(\tilde{x}),y)$, e.g., $l_{1}$ loss for depth estimation. Since the usual prompt space is discrete (text tokens) and challenging to optimize over, we employ the approach introduced in Textual Inversion [22] and instead optimize over the corresponding continuous embedding space. For ease of notation, “prompt space” will implicitly refer to the continuous embedding space instead of the discrete token space. We construct a prompt $c_{w}$ out of $n$ new “placeholder” tokens, i.e., $c_{w}=(c_{w_{1}},\dots,c_{w_{n}})$, and find their corresponding embedding weights $\{w_{i}\}_{i=1}^{n}$ by solving the following optimization problem:
@@
Avoiding $(\tilde{x},y)$ alignment collapse. The adversarial objective in Eq. 1 aims to fool the model $f$. However, it may instead fool the label-conditioning mechanism of the generative model $g$, resulting in $c_{w_{\mathrm{adv}}}$ generating samples $\tilde{x}\sim g(y,c_{w_{\mathrm{adv}}})$ that are not faithful to $y$ (see Fig. 2).
Figure 2: Ways to alleviate the misalignment of the generation with its conditioning. The third column onwards shows some examples of generations from depth maps that do not follow the depth conditioning. See the first and second columns for the original image and its depth label. There are several ways to constrain the generation to alleviate this misalignment. 1. Early stopping involves stopping the adversarial optimization when the loss reaches a certain threshold. The resulting generations from early stopping are shown in the fourth column. 2. SDEdit [51] involves conditioning the generation process on the original image. This mechanism is applied during generation with the adversarial prompts i.e., applying SDEdit to the prompts that generated the images in the third column results in the last column generations. Both SDEdit and early stopping are able to improve the alignment of the generations with depth conditioning.
To avoid this, we further constrain the expressiveness of the generation process. There are several ways to do so.
One way is to use the SDEdit method [50], which conditions the generation process on the original image by starting the denoising process from a noised version of $x$ instead of pure noise. Thus, it constrains the expressive power of the generative model to produce samples closer to the original image $x$.
Additionally, some constraints can be implemented w.r.t $\mathcal{L}_{\mathrm{adv}}$. For the depth estimation task, we employ an early stopping criterion and stop the adversarial optimization when the loss reaches a certain task-specific threshold. For semantic classification, choosing $\mathcal{L}_{\mathrm{adv}}$ to be the negative cross-entropy loss, although natural, may not be a good choice. Indeed, for iWildCam, although we keep the class mask intact, we observed that optimizing the negative cross-entropy loss may lead to the generation of another class somewhere else in the image, e.g. an elephant is generated next to a giraffe, destroying the $(\tilde{x},y)$ alignment. Thus, for iWildCam, we choose to maximize the uncertainty or entropy of the model’s prediction on the generated images. We provide more details in the Appendix Sec. 7.5.2.
Finally, our CLIP [57] guidance loss introduced in Sec. 3.3 further constrains possible perturbations to a target distribution and helps to avoid the generation of non-realistic images.

### v2

@@ Abstract
We present a method to control a text-to-image generative model to produce training data useful for supervised learning. Unlike previous works that employ an open-loop approach via pre-defined prompts to generate new data using either a language model or human expertise, we develop an automated closed-loop system that involves two feedback mechanisms. The first mechanism uses feedback from a given supervised model to find adversarial prompts that result in generated images that maximize the model’s loss and, consequently, expose its vulnerabilities. While these adversarial prompts generate training examples curated for improving the given model, they are not curated for a specific target distribution of interest, which can be inefficient. Therefore, we introduce the second feedback mechanism that can optionally guide the generation process towards a desirable target distribution. We call the method combining these two mechanisms Guided Adversarial Prompts. The proposed closed-loop system allows us to control the training data generation for a given model and target image distribution (see Fig. 1 (Right)). We evaluate on different tasks, datasets, and architectures, with different types of distribution shifts (corruptions, spurious correlations, unseen domains) and illustrate the advantages of the proposed feedback mechanisms compared to open-loop approaches.
@@ 1 Introduction
The quality of data plays a crucial role in training generalizable deep learning models (Taori et al., 2020; Miller et al., 2021; Gadre et al., 2023). For a model to generalize well, its training data should be representative of the test distribution where it will be deployed. However, real-world test conditions change over time, while training datasets are typically collected once and remain static due to high collection costs. This is also in contrast to evidence showing that being able to control the input data is a key contributor to how children are able to learn with only a few examples (Braddick & Atkinson, 2013; Lewis & Maurer, 2005). We, therefore, focus on generating training datasets that can adapt to novel test distributions and are more sample-efficient.
@@
Diffusion generative models (Rombach et al., 2022; Ho et al., 2020; Sohl-Dickstein et al., 2015; Nichol et al., 2022; Saharia et al., 2022) are trained on large-scale collections of images (Schuhmann et al., 2022) and exhibit remarkable generalization abilities by being able to produce realistic images not seen during training. Additionally, unlike static datasets that they are trained on, these generative models allow us to adapt the generation process to produce images that follow a certain conditioning. For example, they can be conditioned on textual prompts (Rombach et al., 2022) or geometric information such as depth maps (Zhang & Agrawala, 2023).
Figure 1: We propose a framework for the automatic generation of training examples curated for a given model and target distribution. Left: The method controls the prompts of a conditional text-to-image generative model via two feedback mechanisms. The ‘model feedback’ ensures that generated examples expose vulnerabilities by fooling a given model, resulting in Adversarial Prompts. The ‘target distribution feedback’ guides the generations towards a desirable target distribution (fog in this example.) Combining both feedback mechanisms results in Guided Adversarial Prompts, leading to generations that both fool the model and fit the desirable target distribution. Right: This approach allows us to find prompts that generate training examples useful for different target distributions, supervised models, and tasks (the oval highlights the effect of defocus blur in the corresponding example.) While the differences in generations for different models might appear subtle or unintuitive, they, indeed, result in more useful training data (see Sec. 4.3).
@@
Recent works explore the use of diffusion models to generate training data for supervised learning with promising results (Sariyildiz et al., 2023; Dunlap et al., 2023; He et al., 2022). They guide the generation process using text prompts to accomplish two goals: 1) produce aligned image-label pairs for supervised training and 2) adapt the generated images to a certain target distribution. However, these methods find conditioning text prompts in an open-loop way, i.e., uninformed of the specific training scenario in hand, by either using a language model (Dunlap et al., 2023) or heuristics (Sariyildiz et al., 2023). In other words, they lack an automatic feedback mechanism that can refine the found text prompts to produce more curated and useful training data.
@@
In this work, we propose two feedback mechanisms to find prompts for generating useful training data. The first mechanism finds prompts that result in generations that maximize the loss of a particular supervised model, thus, reflecting its failure modes. We call them Adversarial Prompts (AP). This mechanism ensures that we find not only novel prompts, which may still produce images that the model already performs well on, but adversarial prompts that produce images with high loss, and, thus, useful for improving the model (Deng et al., 2021) (see exemplar generations for AP in Figs. 1, 3 and 5)
@@
One given model can perform poorly on multiple distribution shifts, and hoping to cover all of the shifts via Adversarial Prompts is an inefficient strategy (e.g., see the difference between the AP generation and the target distribution, fog, in Fig. 1 (Left)). Therefore, we introduce an additional target distribution-informed feedback mechanism that can find prompts that generate images fitting a desirable target distribution we want to adapt to. To implement it, we assume access to a textual description of the target distribution and/or a few unlabeled images from it. We then optimize a similarity metric between CLIP (Radford et al., 2021) embeddings of the generated examples and the target description. We call prompts that combine both mechanisms Guided Adversarial Prompts (GAP). Compare the Adversarial Prompts and Guided Adversarial Prompts in Figs. 1, 3, 4 and 5 to see the effect of the guidance feedback in steering the generations towards a specific target distribution.
@@
We perform evaluations on different tasks (image classification, depth estimation), datasets with distribution shifts (Waterbirds (Sagawa et al., 2019), iWildCam (Beery et al., 2021; Koh et al., 2021), Common Corruptions (Hendrycks & Dietterich, 2019), 3D Common Corruptions (Kar et al., 2022)), and architectures (convolutional and transformer) and show supportive results.
@@ 2 Related Work
Open-loop data generation methods use pre-defined controls to guide the generative process and produce novel training examples. One line of work uses GANs (Jahanian et al., 2021; Chai et al., 2021; Ravuri & Vinyals, 2019) and pre-define a perturbation in their latent space to generate novel examples. More recent works adopt text-to-image diffusion models and use pre-defined prompt templates (Sariyildiz et al., 2023; Yuan et al., 2022; He et al., 2022) or use a language model to generative variations of a given prompt (Yuan et al., 2022). These methods require anticipating the kind of data that will be seen at test-time when defining the prompts. On the other hand, our CLIP guidance mechanism allows us to generate images similar to the target distribution. Dunlap et al. (2023) also approach this problem by using a captioning and language model to summarize a target distribution shift into a text prompt. However, this summarization process is not informed of the generations, and, thus, does not guarantee that the text prompt will guide the generation process to images related to the target distribution. Finally, these methods are not model-informed and do not necessarily generate images useful for training a given model.
@@
Closed-loop data generation methods guide the generation process via an automatic feedback mechanism. They control the latent space of GANs (Besnier et al., 2020) or VAEs (Wong & Kolter, 2020) models, NeRF (Dong et al., 2022), or the space of hand-crafted augmentations (Cubuk et al., 2018) to generate data that maximizes the loss of the network on the generated data. Similarly, Jain et al. (2022) uses an SVM to identify the failure modes of a given model and uses this information to generate training data with a diffusion model. Our method employs a similar adversarial formulation (in conjunction with target distribution guidance) but performs the optimization in the text prompt space of recently developed diffusion models.
@@
“Shallow” data augmentation techniques apply simple hand-crafted transformations to training images to increase data diversity and improve the model’s generalization. Examples of such transformations are color jitter, random crop, and flipping, etc. To produce more diverse augmentations, methods like RandAugment (Cubuk et al., 2020) and AugMix (Hendrycks et al., 2019) combine multiple of such simple transformations, and Mixup (Zhang et al., 2017) and CutMix (Yun et al., 2019) methods use transformations that can combine multiple images. AutoAugment (Cubuk et al., 2018) and adversarial training (Madry et al., 2017) build a closed system to tune the parameters of the applied augmentations but are inherently limited by the expressiveness of the simple transformations. In contrast, our method uses expressive diffusion models, which results in images that are more diverse and realistic than those produced by “shallow” augmentations.
@@
Controlling diffusion models. Methods like ControlNet (Zhang & Agrawala, 2023) and T2I-Adapter (Mou et al., 2023) adapt a pre-trained diffusion model to allow for additional conditioning e.g., edge, segmentation, and depth maps. We employ these models for generation as it allows us to generate paired data for different tasks, given the labels from an existing dataset. Editing methods aim to modify a given image, either via the prompt (Hertz et al., 2022), masks (Couairon et al., 2022), instructions (Brooks et al., 2023) or inversion of the latent space (Mokady et al., 2023; Huberman-Spiegelglas et al., 2023). In contrast, personalization methods aim to adapt diffusion models to a given concept e.g., an object, individual, or style. Popular examples include textual inversion (Gal et al., 2022) and DreamBooth (Ruiz et al., 2023), which aim to find a token to represent a concept given several images of that concept. The former freezes the diffusion model, while the latter fine-tunes it. Extensions of these works learn to represent multiple concepts (Avrahami et al., 2023; Han et al., 2023). In our work, we adopt an approach similar to textual inversion to steer the diffusion model, but our method can also be used with other controlling mechanisms.
@@ 3 Method
Figure 2: Left: An overview of how we generate training data for a given supervised model $f$ and target distribution. Suppose $g$ is generative model that generates images conditioned on a text prompt, $S$ and label, $y$. A supervised model trained to perform a task, e.g., image classification or depth estimation, is denoted by $f$. We aim to find prompts that would generate training data useful for $f$, and we do so via two feedback mechanisms. The first mechanism makes use of feedback from $f$ to get Adversarial Prompts. Since there can be many adversarial examples not relevant to a particular target distribution, we introduce a second feedback mechanism that guides the prompts towards a certain target distribution. Combining the two mechanisms results in Guided Adversarial Prompts. This results in generations that are both relevant to the target distribution and where $f$ does not perform well. Right: Ways to alleviate the misalignment of the generation with its conditioning. The third column onwards shows some examples of generations from depth maps that do not follow the depth conditioning. See the first and second columns for the original image and its depth label. There are several ways to constrain the generation to alleviate this misalignment. 1. Early stopping involves stopping the adversarial optimization when the loss reaches a certain threshold. The resulting generations from early stopping are shown in the fourth column. 2. SDEdit (Meng et al., 2021b) involves conditioning the generation process on the original image. This mechanism is applied during generation with the adversarial prompts i.e., applying SDEdit to the prompts that generated the images in the third column results in the last column generations. Both SDEdit and early stopping are able to improve the alignment of the generations with depth conditioning.
@@ 3.1 Preliminaries
Problem Formulation. We consider the supervised learning problem, where a model $f:\mathcal{X}\to\mathcal{Y}$ learns a mapping from the image space $\mathcal{X}$, to a target space $\mathcal{Y}$, e.g., depth estimation or semantic classification. The model $f$ is trained using a training dataset $\mathcal{D}_{\text{train}}$ and tested on a new set $\mathcal{D}_{\text{test}}$ that exhibits a distribution shift w.r.t. the training data, e.g., corruptions Hendrycks & Dietterich (2019). Our goal is to generate additional synthetic training data $\mathcal{D}_{\text{syn}}$ to adapt the model and improve its performance under the distribution shift.
Text-to-image Diffusion Models. We use the Stable Diffusion (Rombach et al., 2022) text-to-image diffusion model as the basis for our generator $g$. Given a textual prompt $c$, Stable Diffusion is capable of synthesizing realistic images following the textual conditioning. However, in general, for a given task, e.g., depth estimation, a textual prompt alone may not be sufficient for controlling the generation well enough to produce aligned image-label examples.
Generating aligned training examples. We employ the following two approaches to condition the generative model $g$ on a label $y$ from the given training datasets and sample aligned training examples $(\tilde{x},y)$. For the depth estimation task, we use the ControlNet (Zhang & Agrawala, 2023) model which extends the conditioning mechanisms of the Stable Diffusion to accept various spatial modalities, e.g., depth maps and segmentation masks. Specifically, we use ControlNet v1.0 with depth conditioning [^1] https://github.com/lllyasviel/ControlNet. For semantic classification tasks, we utilize the foreground object masks and use an in-painting technique proposed by Lugmayr et al. (2022) that preserves the masked region throughout the denoising process, essentially keeping it intact. These mechanisms provide us with a generative model conditioned both on a text prompt $c$ and label $y$. We denote the resulting distribution modeled by this generative model as $g(y,c)$.
@@ 3.2 Model-Informed Generation with Adversarial Prompt Optimization
Our first feedback mechanism aims at generating training examples that reflect the failure modes of a given model $f$. An automatic way to do so is via adversarial optimization, which finds the “worst case” failure modes of $f$. More precisely, we find a prompt $c$ that generates images $\tilde{x}\sim g(y,c)$ that maximize the supervised loss $\mathcal{L}(f(\tilde{x}),y)$, e.g., $l_{1}$ loss for depth estimation. Since the usual prompt space is discrete (text tokens) and challenging to optimize over, we employ the approach introduced by Gal et al. (2022) and, instead, optimize over the corresponding continuous embedding space. For ease of notation, “prompt space” will implicitly refer to the continuous embedding space instead of the discrete token space. We construct a prompt $c_{w}$ out of $n$ new “placeholder” tokens, i.e., $c_{w}=(c_{w_{1}},\dots,c_{w_{n}})$, and find their corresponding embedding weights $\{w_{i}\}_{i=1}^{n}$ by solving the following optimization problem:
@@
Avoiding $(\tilde{x},y)$ alignment collapse. While the adversarial objective in Eq. 1 aims to fool the model $f$, it may instead fool the label-conditioning mechanism of the generative model $g$, resulting in $c_{w_{\mathrm{adv}}}$ generating samples $\tilde{x}\sim g(y,c_{w_{\mathrm{adv}}})$ that are not faithful to $y$ (see Fig. 2 (Right)). To avoid this collapse, we use the following techniques.
First, we limit the expressivity of the generative model using the SDEdit method (Meng et al., 2021a). It conditions the generation process on the original image $x$ by starting the denoising process from its noised version instead of pure noise, constraining the expressive power of the generative model and producing samples closer to the original image $x$, and, therefore, more faithful to $y$.
Additionally, we implement constraints w.r.t. $\mathcal{L}_{\mathrm{adv}}$. For depth estimation, we employ an early stopping criterion and stop the adversarial optimization when the loss reaches a certain threshold. We also experimented with constraining the embedding weights $w$ to be close to existing vectors from the vocabulary in E.q. 1. However, we found that this significantly reduces the expressivity of the generations. Thus, we do not constrain $w$ during optimization. For image classification, we found that adversarial optimization of the standard cross-entropy loss can lead to the generation of images of another class (e.g., generating elephant, when $y=\mathrm{giraffe}$). We, therefore, maximize the entropy of the model’s predictions, i.e., its uncertainty, instead, which we found to lead to the generation of aligned $(\tilde{x},y)$ examples. Please see Sec. B.5.2 for more details and examples.